# The Wavefront Set: Bounds for the Langlands parameter

### v2

@@ 4.5. Ramified unitary groups
For the ramified unitary groups, the finite group ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})$ on which the cuspidal representation lives is a product of a symplectic and special orthogonal group. For notational purposes, assume ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})=Sp(2n_{1},{\mathbb{F}}_{q})\times SO(n_{2}+n_{2}^{an},n_{2},{\mathbb{F}}_{q})$. We only record the analogue of Propositions 4.5 and 4.7. The proof is simply a combination of the symplectic and orthogonal cases discussed before.
@@ 5. Tempered L-packets of $\mathbf{G_{2}}$
Langlands correspondences for $G_{2}$ were recently constructed in [AX] and [GS1, GS2]. We verify Conjecture 1.6 for $G_{2}$ using the explicit description of L-packets in [AX].
@@ 5.1. The verification of the main conjectures for $G_{2}$
If $\pi$ is not trivial and not generic, then ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)={\mathrm{G}_{2}(a_{1})}.$
@@
By [JLS, §12 and Table $G_{2}$, p. 437], the only possible geometric wavefront sets in $G_{2}$ are ${\mathrm{G}_{2}}$, ${\mathrm{G}_{2}(a_{1})}$, ${A_{1}}$, and 11{1}1. By [LoSa, Theorem 1.2], the minimal orbit ${A_{1}}$ does not occur as a wavefront set. Hence ${\mathrm{G}_{2}}$, ${\mathrm{G}_{2}(a_{1})}$, and 11{1}1 are the only possibilities. The wavefront set of $\pi$ is the regular orbit ${\mathrm{G}_{2}}$ if and only if $\pi$ is generic. On the other hand, the wavefront set is $1$ if and only if $\pi$ is an unramified representation, hence trivial. ∎
@@
Let $\pi$ be a tempered irreducible representation of $G_{2}$. Let ${\mathbb{O}}_{\pi}^{\vee}\in\{{\mathrm{G}_{2}},{\mathrm{G}_{2}(a_{1})},{\widetilde{A}_{1}},{A_{1}},{1}\}$ be the nilpotent part of the Langlands parameter of $\pi$. If ${\mathbb{O}}_{\pi}^{\vee}$ is not the trivial orbit, ${\mathsf{AZ}}(\pi)$ is not generic and $\mathsf{WF}({\mathsf{AZ}}(\pi))={\mathrm{G}_{2}(a_{1})}.$
@@
We fix a maximal ${\mathsf{k}}$-split torus $T$, a Borel subgroup $B$ with $T\subset B$. Let $\Delta=\{\alpha,\beta\}$ be the set of simple roots associated to $B$ where $\alpha$ is a short root. Let $P_{\gamma}=M_{\gamma}U_{\gamma}$, $\gamma\in\Delta\cup\{\emptyset\}$ be the parabolic subgroup associated to $\gamma$. We identify $T$ with ${\mathsf{k}}^{\times}\times{\mathsf{k}}^{\times}$ via $M_{\alpha}\simeq GL_{2}$. We also adopt the notation from [GS1]:
@@
  • $J_{R}(\cdots)$ where $R$ is a parabolic denotes a Langlands quotient of $I_{R}(\cdots)$.
@@
Thanks to the classification of unitary dual of $G_{2}$ by Muic ([Mu]), we can make an explicit calculation of the ${\mathsf{AZ}}$ duals. We also need the following facts:
@@
    If $\pi$ is a nontrivial non-generic irreducible representation of $G_{2}$, $\mathsf{WF}(\pi)$ is ${\mathrm{G}_{2}(a_{1})}$.
We now explain the calculation of the ${\mathsf{AZ}}$ map. Since subquotients of the parabolically induced presentations are preserved by the ${\mathsf{AZ}}$ map, we can consider the ${\mathsf{AZ}}$ map restricted to a finite set of representations with the same cuspidal support. We will show two cases since the rest of calculation is similar: In the following, $=$, $+$ are operations in $K(G_{2})$.
@@
${\mathsf{AZ}}:\{\pi_{\textup{gen}}(\chi)=\pi_{\textup{gen}}(\chi^{-1}),J_{P}(1/2,\delta(\chi)\}\ \Leftrightarrow\ \{J_{P}(1/2,\delta(\chi^{-1}),J_{Q}(1/2,\pi(\chi,\chi^{-1}))\}$
@@
$\pi$ ${\mathbb{O}}_{\pi}$ $d^{\vee}({\mathbb{O}}_{\pi})$ ${\mathsf{AZ}}(\pi)$ $\mathsf{WF}({\mathsf{AZ}}(\pi))$ reference
$I(2,1,1_{{\mathsf{k}}^{\times}},1_{{\mathsf{k}}^{\times}})$ $St_{G_{2}}$ ${\mathrm{G}_{2}}$ 11{1}1 $1_{G_{2}}$ 11{1}1 Prop. 3.1 (ii)-(iii) Prop. 3.2 (ii)-(iii) [GS1]
@@
$1_{G_{2}}$ 11{1}1 ${\mathrm{G}_{2}}$ $St_{G_{2}}$ ${\mathrm{G}_{2}}$
$I(1,0,1_{{\mathsf{k}}^{\times}},1_{{\mathsf{k}}^{\times}})$ $\pi_{\textup{gen}}(1)$ ${\mathrm{G}_{2}(a_{1})}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(1,\pi(1,1)$ ${\mathrm{G}_{2}(a_{1})}$
$\pi_{\textup{deg}}^{r}(1)$ ${\mathrm{G}_{2}(a_{1})}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{P}(\frac{1}{2},\delta(1))$ ${\mathrm{G}_{2}(a_{1})}$
@@
$J_{P}(\frac{1}{2},\delta(1))$ ${A_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $\pi_{\textup{deg}}^{r}(1)$ ${\mathrm{G}_{2}(a_{1})}$
$J_{Q}(1,\pi(1,1))$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi_{\textup{gen}}(1)$ ${\mathrm{G}_{2}}$
@@
$J_{Q}(\frac{1}{2},\pi(1,\chi))$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi_{\textup{gen}}(\chi)$ ${\mathrm{G}_{2}}$
@@
$J_{Q}(1,\pi(\chi,\chi^{-1}))$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi_{\textup{gen}}(\chi)$ ${\mathrm{G}_{2}}$
$i_{B}^{G}\chi$ irred. $i_{B}^{G}\chi$ 11{1}1 ${\mathrm{G}_{2}}$ $i_{B}^{G}\chi$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{P}(\frac{1}{2},\tau)\\ \tau\simeq\tau^{\vee},\ \omega_{\tau}=1\end{array}$ $\delta_{P}(\tau)$ ${A_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{P}(\frac{1}{2},\tau)$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.1-(i)]
$J_{P}(\frac{1}{2},\tau)$ 11{1}1 ${\mathrm{G}_{2}}$ $\delta_{P}(\tau)$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{P}(0,\tau),\tau\simeq\tau^{\vee}\\ \omega_{\tau}\neq 1\end{array}$ $I_{P}(\tau)_{\textup{gen}}$ 11{1}1 ${\mathrm{G}_{2}}$ $I_{P}(\tau)_{\textup{deg}}$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.1-(ii)]
$I_{P}(\tau)_{\textup{deg}}$ 11{1}1 ${\mathrm{G}_{2}}$ $I_{P}(\tau)_{\textup{gen}}$ ${\mathrm{G}_{2}}$
$i_{P}^{G}\tau$ irred. $i_{Q}^{G}\tau$ 11{1}1 ${\mathrm{G}_{2}}$ $i_{P}^{G}\tau$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{Q}(\frac{1}{2},\tau)\\ \tau\simeq\tau^{\vee},\ \omega_{\tau}=1\end{array}$ $\delta_{Q}(\tau)$ ${\widetilde{A}_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(\frac{1}{2},\tau)$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.2-(i)]
$J_{Q}(\frac{1}{2},\tau)$ 11{1}1 ${\mathrm{G}_{2}}$ $\delta_{Q}(\tau)$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{Q}(1,\tau),\ \tau\simeq\tau^{\vee}\\ \tau\textrm{ dihedral},\ \omega_{\tau}\neq 1\end{array}$ $\pi_{\textup{gen}}(\tau)$ ${\mathrm{G}_{2}(a_{1})}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(1,\tau)$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.2-(ii)]
$J_{Q}(1,\tau)$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi_{\textup{gen}}(\tau)$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{Q}(0,\tau),\tau\simeq\tau^{\vee}\\ \omega_{\tau}\neq 1\end{array}$ $I_{Q}(\tau)_{\textup{gen}}$ 11{1}1 ${\mathrm{G}_{2}}$ $I_{Q}(\tau)_{\textup{deg}}$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.2-(iii)]
$I_{Q}(\tau)_{\textup{deg}}$ 11{1}1 ${\mathrm{G}_{2}}$ $I_{Q}(\tau)_{\textup{gen}}$ ${\mathrm{G}_{2}}$
$i_{Q}^{G}\tau$ irred. $i_{Q}^{G}\tau$ 11{1}1 ${\mathrm{G}_{2}}$ $i_{Q}^{G}\tau$ ${\mathrm{G}_{2}}$
$\pi$, nonsingular s.c. $\pi$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi$ ${\mathrm{G}_{2}(a_{1})}$ [Kal]
or ${\mathrm{G}_{2}}$
$\pi$, singular s.c. $\pi$ $\begin{array}[]{c}{\widetilde{A}_{1}},\ {A_{1}}\\ \textrm{or }{\mathrm{G}_{2}(a_{1})}\end{array}$ ${\mathrm{G}_{2}(a_{1})}$ $\pi$ ${\mathrm{G}_{2}(a_{1})}$
Table 1. Irrreducible tempered representations of $G_{2}$ and their ${\mathsf{AZ}}$-duals
@@
Conjecture 1.9 holds for any irreducible tempered representation $\pi$ on $G_{2}$.
@@
This is immediate from the previous results. ∎
5.2. Depth-zero representations

### v3

@@ 4.6. Ramified unitary groups
For the ramified unitary groups, the finite group ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})$ on which the cuspidal representation lives is a product of a symplectic and special orthogonal group. For notational purposes, assume ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})=Sp(2n_{1},{\mathbb{F}}_{q})\times SO(n_{2}+n_{2}^{an},n_{2},{\mathbb{F}}_{q})$. We only record the analogue of Propositions 4.6 and 4.10. The proof is simply a combination of the symplectic and orthogonal cases discussed before.
@@ 5. Tempered $L$-packets of ${\mathbf{G}}_{2}$}
Langlands correspondences for $G_{2}$ were recently constructed in [AX] and [GS1, GS2] when the characteristic of $\mathsf{k}$} is zero. We expect that their constructions coincide.
Here, we prove Conjecture 1.9 using the explicit description of $L$-packets in [GS1, GS2] when $\mathsf{k}$} has characteristic 0. We also compute the unramified wavefront sets of depth zero tempered representations of $G_{2}$ using the description of Langlands parameters in [AX].
@@ 5.1. The verification of the main conjectures for $G_{2}$
If $\pi\in\mathsf{Irr}\,G_{2}$ is a nontrivial non-generic, then $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)={\mathrm{G}_{2}(a_{1})}.$
@@
By [JLS, §12 and Table $G_{2}$, p. 437], the only possible geometric wavefront sets in $G_{2}$ are $\mathrm{G}_{2}$}, $\mathrm{G}_{2}(a_{1})$}, $A_{1}$}, and 1{1}. By [LoSa, Theorem 1.2], the minimal orbit $A_{1}$} does not occur as a wavefront set. Hence $\mathrm{G}_{2}$}, $\mathrm{G}_{2}(a_{1})$}, and 1{1} are the only possibilities. The wavefront set of $\pi$ is the regular orbit $\mathrm{G}_{2}$} if and only if $\pi$ is generic. On the other hand, the wavefront set is $1$ if and only if $\pi$ is an unramified character, hence trivial. ∎
@@
Let $\pi$ be a tempered irreducible representation of $G_{2}$. Let ${\mathbb{O}}_{\pi}^{\vee}\in\left\{{\mathrm{G}_{2}},{\mathrm{G}_{2}(a_{1})},{\widetilde{A}_{1}},{A_{1}},{1}\right\}$ be the nilpotent part of the Langlands parameter of $\pi$.
  1. (1)
    Suppose $\pi\neq\mathsf{St}_{G_{2}}$. If ${\mathbb{O}}_{\pi}^{\vee}$ is not the trivial orbit, ${\mathsf{AZ}}(\pi)$ is not generic and $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))={\mathrm{G}_{2}(a_{1})}.$
  2. (2)
    If $\pi=\mathsf{St}_{G_{2}}$, we have ${\mathbb{O}}^{\vee}_{\pi}={\mathrm{G}_{2}}$, ${\mathsf{AZ}}(\pi)=1_{G_{2}}$ and $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=1$.
  3. (3)
    If ${\mathbb{O}}^{\vee}_{\pi}$ is trivial, $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))$ is either $\mathrm{G}_{2}$} or $\mathrm{G}_{2}(a_{1})$}.
@@
(2) and (3) are straightforward.
Now suppose $\pi\neq\mathsf{St}_{G_{2}}$ and ${\mathbb{O}}_{\pi}^{\vee}\neq 1$. We fix a maximal $\mathsf{k}$}-split torus $T$, a Borel subgroup $B$ with $T\subset B$. Let $\Delta=\left\{\alpha,\beta\right\}$ be the set of simple roots associated to $B$ where $\alpha$ is a short root. Let $P_{\gamma}=M_{\gamma}U_{\gamma}$, $\gamma\in\Delta\cup\left\{\emptyset\right\}$ be the parabolic subgroup associated to $\gamma$. We identify $T$ with ${\mathsf{k}}^{\times}\times{\mathsf{k}}^{\times}$ via $M_{\alpha}\simeq GL_{2}$. We also adopt the notation from [GS1]:
@@
  • $J_{R}(\cdots)$, where $R$ is a parabolic subgroup, denotes a Langlands quotient of $I_{R}(\cdots)$.
@@
To calculate $\mathsf{AZ}$} duals explicitly, we use various calculations in the classification of unitary dual of $G_{2}$ by Muic ([Mu]). These are also summarized in [GS1, Propositions 3.1, 3.2]. We also need the following facts for the calculation:
@@
    Lemma 5.1.
We now explain the calculation of the $\mathsf{AZ}$} map. Since subquotients of the parabolically induced presentations are preserved by the $\mathsf{AZ}$} map, we can consider the $\mathsf{AZ}$} map restricted to a finite set of representations with the same cuspidal support. We will show two cases since the rest of calculation is similar: In the following, ==, ++ are operations in $K(G_{2})$.
@@
${\mathsf{AZ}}:\left\{\pi_{\textup{gen}}(\chi)=\pi_{\textup{gen}}(\chi^{-1}),J_{P}(1/2,\delta(\chi)\right\}\ \Leftrightarrow\ \left\{J_{P}(1/2,\delta(\chi^{-1}),J_{Q}(1/2,\pi(\chi,\chi^{-1}))\right\}$
@@
The above calculation is summarized in Table 1. The calculation of ${\mathbb{O}}_{\pi}^{\vee}$ in the table is from [GS1, §3.5]. The proof of the theorem is now complete. ∎
$\pi$ ${\mathbb{O}}_{\pi}^{\vee}$ $d^{\vee}({\mathbb{O}}_{\pi}^{\vee})$ ${\mathsf{AZ}}(\pi)$ $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))$ reference in [GS1]
$I(2,1,1_{{\mathsf{k}}^{\times}},1_{{\mathsf{k}}^{\times}})$ $\mathsf{St}_{G_{2}}$ $\mathrm{G}_{2}$} 1{1} $_{G_{2}} 1{1}$
Prop. 3.1 (ii)-(iii)
Prop. 3.2 (ii)-(iii)
@@
$_{G_{2}} 1{1}$ $\mathrm{G}_{2}$} $\mathsf{St}_{G_{2}}$ $\mathrm{G}_{2}$}
$I(1,0,1_{{\mathsf{k}}^{\times}},1_{{\mathsf{k}}^{\times}})$ $\pi_{\textup{gen}}(1)$ $\mathrm{G}_{2}(a_{1})$} $\mathrm{G}_{2}(a_{1})$} $J_{Q}(1,\pi(1,1))$ $\mathrm{G}_{2}(a_{1})$}
$\pi_{\textup{deg}}(1)$ $\mathrm{G}_{2}(a_{1})$} $\mathrm{G}_{2}(a_{1})$} $J_{P}(\frac{1}{2},\delta(1))$ $\mathrm{G}_{2}(a_{1})$}
@@
$J_{P}(\frac{1}{2},\delta(1))$ $A_{1}$} $\mathrm{G}_{2}(a_{1})$} $\pi_{\textup{deg}}(1)$ $\mathrm{G}_{2}(a_{1})$}
$J_{Q}(1,\pi(1,1))$ 1{1} $\mathrm{G}_{2}$} $\pi_{\textup{gen}}(1)$ $\mathrm{G}_{2}$}
@@
$J_{Q}(\frac{1}{2},\pi(1,\chi))$ 1{1} $\mathrm{G}_{2}$} $\pi_{\textup{gen}}(\chi)$ $\mathrm{G}_{2}$}
@@
$J_{Q}(1,\pi(\chi,\chi^{-1}))$ 1{1} $\mathrm{G}_{2}$} $\pi_{\textup{gen}}(\chi)$ $\mathrm{G}_{2}$}
$i_{B}^{G}\chi$ irred. $i_{B}^{G}\chi$ 1{1} $\mathrm{G}_{2}$} $i_{B}^{G}\chi$ $\mathrm{G}_{2}$}
$\begin{array}[]{c}I_{P}(\frac{1}{2},\tau)\\ \tau\simeq\tau^{\vee},\ \omega_{\tau}=1\end{array}$ $\delta_{P}(\tau)$ $A_{1}$} $\mathrm{G}_{2}(a_{1})$} $J_{P}(\frac{1}{2},\tau)$ $\mathrm{G}_{2}(a_{1})$} Prop. 3.1-(i)
$J_{P}(\frac{1}{2},\tau)$ 1{1} $\mathrm{G}_{2}$} $\delta_{P}(\tau)$ $\mathrm{G}_{2}$}
$\begin{array}[]{c}I_{P}(0,\tau),\tau\simeq\tau^{\vee}\\ \omega_{\tau}\neq 1\end{array}$ $I_{P}(\tau)_{\textup{gen}}$ 1{1} $\mathrm{G}_{2}$} $I_{P}(\tau)_{\textup{deg}}$ $\mathrm{G}_{2}(a_{1})$} Prop. 3.1-(ii)
$I_{P}(\tau)_{\textup{deg}}$ 1{1} $\mathrm{G}_{2}$} $I_{P}(\tau)_{\textup{gen}}$ $\mathrm{G}_{2}$}
$i_{P}^{G}\tau$ irred. $i_{P}^{G}\tau$ 1{1} $\mathrm{G}_{2}$} $i_{P}^{G}\tau$ $\mathrm{G}_{2}$}
$\begin{array}[]{c}I_{Q}(\frac{1}{2},\tau)\\ \tau\simeq\tau^{\vee},\ \omega_{\tau}=1\end{array}$ $\delta_{Q}(\tau)$ $\widetilde{A}_{1}$} $\mathrm{G}_{2}(a_{1})$} $J_{Q}(\frac{1}{2},\tau)$ $\mathrm{G}_{2}(a_{1})$} Prop. 3.2-(i)
$J_{Q}(\frac{1}{2},\tau)$ 1{1} $\mathrm{G}_{2}$} $\delta_{Q}(\tau)$ $\mathrm{G}_{2}$}
$\begin{array}[]{c}I_{Q}(1,\tau),\ \tau\simeq\tau^{\vee}\\ \tau\textrm{ dihedral},\ \omega_{\tau}\neq 1\end{array}$ $\pi_{\textup{gen}}(\tau)$ $\mathrm{G}_{2}(a_{1})$} $\mathrm{G}_{2}(a_{1})$} $J_{Q}(1,\tau)$ $\mathrm{G}_{2}(a_{1})$} Prop. 3.2-(ii)
$J_{Q}(1,\tau)$ 1{1} $\mathrm{G}_{2}$} $\pi_{\textup{gen}}(\tau)$ $\mathrm{G}_{2}$}
$\begin{array}[]{c}I_{Q}(0,\tau),\tau\simeq\tau^{\vee}\\ \omega_{\tau}\neq 1\end{array}$ $I_{Q}(\tau)_{\textup{gen}}$ 1{1} $\mathrm{G}_{2}$} $I_{Q}(\tau)_{\textup{deg}}$ $\mathrm{G}_{2}(a_{1})$} Prop. 3.2-(iii)
$I_{Q}(\tau)_{\textup{deg}}$ 1{1} $\mathrm{G}_{2}$} $I_{Q}(\tau)_{\textup{gen}}$ $\mathrm{G}_{2}$}
$i_{Q}^{G}\tau$ irred. $i_{Q}^{G}\tau$ 1{1} $\mathrm{G}_{2}$} $i_{Q}^{G}\tau$ $\mathrm{G}_{2}$}
$\pi$, unipotent s.c. $\pi$ $\mathrm{G}_{2}(a_{1})$} $\mathrm{G}_{2}(a_{1})$} $\pi$ $\mathrm{G}_{2}(a_{1})$} §3.5-(2)
$\pi$, singular s.c. $\pi$ ${\widetilde{A}_{1}},\ {A_{1}}$ $\mathrm{G}_{2}(a_{1})$} $\pi$ $\mathrm{G}_{2}(a_{1})$} §3.5-(3),(4)
$\pi$, nonsingular s.c. $\pi$ 1{1} $\mathrm{G}_{2}$} $\pi$ $\mathrm{G}_{2}(a_{1})$} or $\mathrm{G}_{2}$} §3.5-(5)
Table 1. Irreducible tempered representations of $G_{2}$ and their $\mathsf{AZ}$}-duals
@@
Conjecture 1.9-(1) holds for any irreducible tempered representation $\pi$ on $G_{2}$.
@@
Conjecture 1.9-(1) is immediate from Table 1. ∎
Corollary 5.4.
For the $L$-packets in [GS1, GS2], Conjecture 1.9-(2) holds.
Proof.
Write $\varphi:WD_{\mathsf{k}}\longrightarrow{}^{L}\!G$ for a Langlands parameter where $WD_{\mathsf{k}}=W_{\mathsf{k}}\times SL_{2}({\mathbb{C}})$. Conjecture 1.9-(2) results from case by case checking summarized in Table 2. In the following table, ${\mathbb{O}}_{\varphi}^{\vee}$ and $\Pi_{\varphi}$ denote the nilpotent orbit and the $L$-packet associated to $\varphi$ respectively.
${\mathbb{O}}_{\varphi}$ $d^{\vee}({\mathbb{O}}_{\varphi}^{\vee})$ $\varphi(W_{\mathsf{k}})$ $\Pi_{\varphi}$ ${\mathsf{AZ}}(\Pi_{\varphi})$ $\pi$: $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))$ reference on $\varphi$
$d^{\vee}({\mathbb{O}}_{\varphi}^{\vee})$ in [GS1]
$\mathrm{G}_{2}$} 1 1 $\left\{\mathsf{St}_{G_{2}}\right\}$ $1_{G_{2}}$ $\Pi_{\varphi}$ §3.5-(1)
$\mathrm{G}_{2}(a_{1})$} $\mathrm{G}_{2}(a_{1})$} 1 $\left\{\pi_{\textup{gen}}(1),\pi_{\textup{deg}}(1),\pi(1)\right\}$ $\begin{array}[]{l}\{\ J_{Q}(1,\pi(1,1))\\ \quad J_{P}(\frac{1}{2},\delta(1)),\pi(1)\ \}\end{array}$ $\Pi_{\varphi}$ §3.5-(2)
$\mu_{2}$ $\begin{array}[]{c}\{\ \pi_{\textup{gen}}(\chi)\textrm{ with }\chi^{2}=1,\\ \quad\pi(-1)\textrm{ unipotent s.c.}\ \}\end{array}$ $\left\{J_{Q}(\frac{1}{2},\pi(1,\chi)),\pi(-1)\right\}$ $\Pi_{\varphi}$
$\mu_{3}$ $\begin{array}[]{l}\{\ \pi_{\textup{gen}}(\chi)\textrm{ with }\chi^{3}=1,\\ \pi(\omega),\pi(\omega^{2}):\textrm{ unip. s.c.}\ \}\end{array}$ $\begin{array}[]{l}\{\ J_{Q}(\frac{1}{2},\pi(1,\chi)),\\ \quad\pi(\omega),\pi(\omega^{2})\ \}\end{array}$ $\Pi_{\varphi}$
$S_{3}$ $\pi_{\textup{gen}}(\tau)$ $J_{Q}(1,\tau)$ $\Pi_{\varphi}$
$\widetilde{A}_{1}$} $\mathrm{G}_{2}(a_{1})$} $\left\{\delta_{P}(\tau),\pi(-1)\right\}$ $\left\{J_{P}(\frac{1}{2},\tau),\pi(-1)\right\}$ $\Pi_{\varphi}$ §3.5-(3)
$\pi(-1)$ singular s.c.
$A_{1}$} $\mathrm{G}_{2}(a_{1})$} $\left\{\delta_{Q}(\tau),\pi(-1)\right\}$ $\left\{J_{P}(\frac{1}{2},\tau),\pi(-1)\right\}$ $\Pi_{\varphi}$ §3.5-(4)
$\pi(-1)$ singular s.c.
11 $\mathrm{G}_{2}$} $\left\{i_{B}^{G}\chi,\textrm{ irred}\right\}$, $\Pi_{\varphi}$ $\begin{array}[]{l}\textrm{generic}\\ \textrm{ element}\\ \ \textrm{\ in }\Pi_{\varphi}\end{array}$ Prop 3.3
$\left\{i_{P}^{G}\tau\right\}$ or $\left\{i_{Q}^{G}\tau\right\}$ §3.5, p17
s.c. $L$-pacekt §3.5-(5)
Table 2. Tempered $L$-packets of $G_{2}$
5.2. Unramified wavefront sets of depth-zero representations